2. Sketch and find the equation of the curve. x = t2 – 3, y = t2 + 2, -2 st s2
Answer #1

Hello

x=+²- 30 y=+²+2-0 -astea then ost²<4 Date 1 - 3 st²-3&1 zst+256 T-250sl a <yeol 0.- € we get x - y = -3-2 1 y=x+5 1. Sketch i

Thanks